Lord have mercy heres another one it says to solve by elimination method
0.3x-0.2y=4
0.2x+0.5y=-75/17
Pleeese someone help. Thanks so much

Answer by edjones(8007)

A) 0.3x-0.2y=4
B) 0.2x+0.5y=-75/17
.
A) 3x-2y=40 multiply by 10
B) 2x+5y=-750/17
B) 34x+85y=-750 Multiply by 17
.
B) 68x+170y=-1500
A) 255x-170y=3400
------------------------------add
=>323x=1900
.
x=100/17
.
B)
.2(100/17)+.5y=-75/17
20/17+.5y=-75/17
.5y=-95/17
y=-190/17
.
Ed